If Kevin invests $50,000 into an account that earns 6.15% interest compounded monthly, how long will it take his money to double? (round to one decimal place)
A.135.6 years
B.11.6 years
C.11.3 years
D.9 years

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Rule of 70 would exclude a and d. 70/6.15=11.38 favoring c
P=Po(1+(0.0615/12)^nt
2=(1.005125)^nt
ln2=nt ln 1.005125
nt=135.6 compounding times
divide by 12 to get years , which is 11.3
C

Answer by ikleyn(52875)

.

In such problems,  the quantum of the time is   a MONTH.


Therefore,  would the problem be formulated  ACCURATELY  by its composer,
it should ask about the  MONTHS --- not about the decimal parts of the year.


DIAGNOSIS:   The problem formulation is   lame.


/\/\/\/\/\/\/\/

For the composer's info :

    A year is 12 months - - - not 10 months.


One more notice specially related to the solution by @Boreal :

    nt=135.6 compounding times  CAN NOT be non-integer value.

    It is NECESSARY an INTEGER.   It MUST be an integer.


So, the correct answer in this case is nt = 136 compounding times,
which is 136 months divided by 12 months = 11 years and 4 months.


        *******************************************
            DO NOT ACCEPT any different answer.
        *******************************************